(2)(10pts) A ship has primary navigation system X and backup system Y. X is the event X works and Y is the event Y works.
Answer #1


-out come probability - Subexperiment 11 eller PROQ P(x) = 1- P(x) = 1-q... P (XC) = 9. P(x(x) = p. P(11x)1-P(Y|x) = 3-5P(x) = P(+/+) » +P (rl xe) - ptt. P(rc) = 1- P(7) - gtp-t. (1) P ( xnx) = P(x/x.). P(x) = P(Y/x). (1-P (XC) P(xnyc) - P(10x)P(x1) = P(x0Y) P(4) = P(x-2) (P++) (from previous) P(xUY) = 3-P(xuyje - 1-P (renre) (by De Morgan - 1- ((1-)a) Morgans = P (